OpenStudy (anonymous):

The sum of two numbers is 10. Three times the larger decreased by twice the smaller is 15. What are the numbers

OpenStudy (anonymous):

3 and 7 i think

OpenStudy (anonymous):

a + b = 10 3a - 2b = 15

OpenStudy (anonymous):

7*3=21 3*2=6 21-6=15

OpenStudy (anonymous):